# Generalized Harnack Inequality for Mean Curvature Flow and Ancient Solutions

### v1

@@ 1 Introduction
Let $M^{d}$ be a smooth $d$-dimensional manifold. A smooth one-parameter family of hypersurfaces $\{M_{t}^{d}\}_{t\in I}$ is a smooth solution to mean curvature flow iff there exists a smooth one parameter family of immersions $X_{t}:M^{d}\to\mathbb{R}^{d+1}$ which satisfies the equation
@@
A standard method of analyzing singularities is to use blowup arguments. As the solution approaches singular time, we ’magnify’ at larger and larger scale near the point where singularities form, and take ’the limit’ of the rescaled solutions. In the end, one is left with a solution to the mean curvature flow which is defined for all negative time. Such solutions are called ancient solutions. Therefore, the study of ancient solutions is important since they model how singularities form.
@@
Corollary 1.1 (Recovery of classical Harnack inequality)
@@
In the 2nd half of the paper, we consider ancient solutions to mean curvature flow. In [2], Huisken and Sinestrari provided several characterizations of shrinking spheres for closed, convex ancient solution to mean curvature flow.
Theorem 1.2 (cf. [2], theorem 1.1)
@@
(2) $\liminf_{t\to-\infty}\min_{M^{d}\times\{t\}}\frac{\lambda_{1}}{H}>0$ i.e $M_{t}^{d}$ is uniformly strictly convex.
(3) $\limsup_{t\to-\infty}\frac{diam(M_{t}^{d})}{\sqrt{-t}}<\infty$.
@@
i.e $M_{t}^{d}$ is uniformly strictly convex.
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}Hdiam(M^{d}_{t})<\infty$
where $diam(M^{d}_{t})$ denote the diameter of $M_{t}^{d}\subset\mathbb{R}^{d+1}$.
@@
for either $k=1$ or $k=d-1$. We also impose a ’flatness at low curvature region’ condition, that is for every $\epsilon>0$, there exists $\delta>0$ so that whenever
@@ 2 Generalized Harnack inequality
Proof of theorem 1.1 : For simplicity, we define
@@
$Z_{\epsilon,\hat{T}}(V)=(A+\epsilon Hg)(V,V)+\frac{2}{1+\epsilon d}\nabla H\cdot V+\frac{H}{2t}+\frac{1}{(1+\epsilon d)^{2}}\nabla_{t}H+\frac{\epsilon}{(1+\epsilon d)^{2}}H^{3}+\frac{3\epsilon}{4(1+\epsilon d)}\phi(\hat{T})^{2}H$
@@
for all $V\in TM^{d}$, $0<t\leq\hat{T}<T$. Then theorem 1.1 follows by first letting $\epsilon\to\epsilon_{0}$, then $t=\hat{T}\in(0,T)$. Since $(M^{d},X_{t})_{t\in[0,T)}$ is mean convex solution of mean curvature flow, by the strong maximum principle either $H>0$ for all $t\in(0,T)$ or $H=0$. The theorem trivially holds when $H=0$, thus WLOG we assume $H>0$ for $t>0$. Then for any $\epsilon>\epsilon_{0}$,
@@
For each $(p_{0},t_{0})\in M^{d}\times(0,\hat{T}]$, we let $(U,\{x_{a}\})$ to be a geodesic normal coordinates w.r.t $g(t_{0})$ centered at $p_{0}$ so that the coordinate vectors at $p_{0}$ are the eigenvectors of the 2nd fundamental form i.e at $(p_{0},t_{0})$,
@@
$+(-4\nabla_{c}A_{ab}V_{b}-4\epsilon\nabla_{c}Hg_{ab}V_{b}-4\delta\nabla_{t}A_{ac}+4\delta HA_{ab}A_{bc})\nabla_{c}V_{a}+4\delta H\nabla_{a}A_{bc}A_{bc}V_{a}+2\delta A_{ab}A_{bc}\nabla_{c}HV_{a}$
$+4\delta^{2}HA_{ab}\nabla_{t}A_{ab}+2\delta^{2}A_{ab}\nabla_{a}H\nabla_{b}H-2\delta^{2}H^{2}A_{ab}A_{bc}A_{ca}+2\epsilon\delta^{2}|A|^{2}H^{3}+g^{\prime}(t)H-6\epsilon\delta^{2}H|\nabla H|^{2}$
We plug in our choice of local extension, and compute each terms separately.
@@
Combine all the terms, then we obtain
@@
$(\nabla_{t}-\Delta)Z_{\epsilon,\hat{T}}(V)=|A|^{2}Z_{\epsilon,\hat{T}}(V)-\frac{4f}{\delta}[(A_{ab}+\epsilon Hg_{ab})V_{a}V_{b}+2\delta\nabla_{a}HV_{a}+\delta^{2}\nabla_{t}H]+(6\epsilon^{2}\delta H^{2}-4\epsilon f)\nabla_{a}HV_{a}$
$+g^{\prime}(t)H-2f^{2}H-4f\epsilon\delta H^{3}-2\epsilon df^{2}H$
$=|A|^{2}Z_{\epsilon,\hat{T}}(V)-\frac{4f}{\delta}[Z_{\epsilon,\hat{T}}(V)-g(t)H-\epsilon\delta^{2}H^{3}]+(6\epsilon^{2}\delta H^{2}-4\epsilon f)\nabla_{a}HV_{a}+g^{\prime}(t)H-2f^{2}H-4f\epsilon\delta H^{3}-2\epsilon df^{2}H$
@@
and $(p_{0},t_{0})\in M^{d}\times(0,\hat{T}]$ and $\phi(\hat{T})=\sup_{M^{d}\times(0,\hat{T}]}H$. One can directly check that above two inequalities indeed hold. Thus proving the desired differential inequality.
@@
for $V\in TM^{d}$, $t\in(0,\hat{T}]$, hence prove theorem 1.1.
@@
Corollary 2.1 (Harnack inequality for ancient solutions)
@@
Proof of Corollary 2.1 : For each fixed $V\in TM^{d}$, $t\in(-\infty,0)$, let $-\alpha<t<0$ and consider $\{M_{s}^{d}\}_{s\in[-\alpha,0)}$. Since $\phi(t)\geq\sup_{M^{d}\times(-\alpha,t]}H$, theorem 1.1 implies
@@
for all $V\in TM^{d}$, $t\in(-\alpha,0)$. Thus letting $\alpha\to\infty$, we obtain corollary 2.1.
@@
Proof: Let $\gamma:[t_{1},t_{2}]\to M^{d}$ be any smooth curve connecting $p_{1}$ and $p_{2}$. Dividing the Harnack inequality by $H$ and setting $V=\frac{1}{2(1+\epsilon_{0}d)}\gamma^{\prime}(t)$, we have
@@
thus proving corollary 2.2.
@@ 3 Ancient solutions
$\limsup_{t\to-\infty}\phi(t)diam(M_{t}^{d})<\infty\iff\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}Hdiam(M_{t}^{d})<\infty$
Proof : By definition of $\phi$,
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}Hdiam(M_{t}^{d})\leq\limsup_{t\to-\infty}\phi(t)diam(M_{t}^{d})$
@@
To prove the reverse direction, note that $t\to diam(M_{t}^{d})=\max_{x,y\in M^{d}}|X_{t}(x)-X_{t}(y)|$ is nonincreasing in $t$. We give a brief sketch of why this is true. Since we are working with smooth solutions to the mean curvature flow, for each $t\in(-\infty,0)$
$diam(M_{t}^{d})=\max_{x,y\in M^{d}}|X_{t}(x)-X_{t}(y)|>0$
@@
$\frac{d}{dt}diam(M_{t}^{d})=\frac{\partial}{\partial t}|X_{t}(p)-X_{t}(q)|=\langle\frac{X_{t}(p)-X_{t}(q)}{diam(M_{t}^{d})},\vec{H}(p,t)-\vec{H}(q,t)\rangle$
@@
$0\geq\langle\frac{X_{t}(p)-X_{t}(q)}{diam(M_{t}^{d})},\vec{H}(p,t)-\vec{H}(q,t)\rangle+\frac{2d}{diam(M_{t}^{d})}=\frac{d}{dt}diam(M_{t}^{d})+\frac{2d}{diam(M_{t}^{d})}$
which proves that $diam(M_{t}^{d})$ is nonincreasing in $t$.
@@
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}Hdiam(M_{t}^{d})\leq C<\infty$
@@
Since $diam(M_{t}^{d})\leq diam(M_{s}^{d})$, we have
$(\phi(t)-\epsilon)diam(M_{t}^{d})\leq\max_{M^{d}\times\{s\}}Hdiam(M_{s}^{d})\leq C+1$
@@
$\limsup_{t\to-\infty}\phi(t)diam(M^{d}_{t})\leq C+1<\infty$
thus proving the lemma.
@@
Proof : Define a tensor
@@
$\lambda_{i}\lambda_{d}=sec(e_{i},e_{d})=0$
hence $\lambda_{i}=0$. Here, $sec(e_{i},e_{d})$ is the sectional curvature in directions $e_{i}$ and $e_{d}$. This together with the fact that $E(t)$ has rank $l>0$ implies that at each $(p,t)\in M^{d}\times(-\delta,0)$, the eigenvalues of $A_{(p,t)}$ are either 00 or $\lambda_{d}$, with the corresponding eigenspaces having dimension $d-l$ and $l$ respectively. Thus for any possible $l>0$, $A\geq 0$ in $M^{d}\times(-\delta,0)$. Then by continuity, $A\geq 0$ up to time $t=0$ as well, thus proving the first assertion.
@@
for any $q\in M^{d}$.
Proof of theorem 1.3 : Trivially $(1)$ implies $(2)-(5)$. By [2], theorem 1.1, $(2)$ implies $(1)$. We claim that either $(4)$ or $(5)$ implies $(3)$. By the weak maximum principle, it is easy to see that for all sufficiently small $t$
@@
$diam(M_{t}^{d})\leq 2C_{2}\sqrt{-t}+diam(M_{T_{0}}^{d})\leq C_{3}\sqrt{-t}\leq\frac{C_{4}}{\phi(t)}$
@@
$\limsup_{t\to-\infty}\phi(t)diam(M_{t}^{d})<\infty$
@@
is the 1st eigenvalue of $Hg-A$, we conclude that
@@
Then lemma 3.2 implies that $M_{\infty}^{d}$ is closed and $X_{\infty}(\cdot,0)$ is an isometric embedding into $\mathbb{R}^{d+1}$ with its image being a round sphere. In particular, the local smooth convergence at $t=0$ is actually global smooth convergence, hence the rescaled 2nd fundamental form $\phi(t_{j})^{-1}A(\cdot,t_{j})$ converge uniformly smoothly to the 2nd fundamental form of a round sphere of fixed radius as $j\to\infty$. Since a sphere is uniformly strictly convex, there exists $j_{0}$ so that for all $j\geq j_{0}$
@@
$\limsup_{t\to-\infty}\frac{diam(M_{t}^{d})}{\sqrt{-t}}\leq C\limsup_{t\to-\infty}\phi(t)diam(M_{t}^{d})<\infty$
thus by [2], theorem 1.1, $\{M_{t}^{d}\}_{t\in(-\infty,0)}$ is a family of shrinking round spheres. Therefore either (3), (4) or (5) implies (1), thus completing the proof of theorem 1.3.
Proof of theorem 1.4 : Assume otherwise. Then
@@
defined for $t\in(-\infty,0]$. Then as in the proof of theorem 1.3, we can extract a limit flow $X_{\infty}:M^{d}_{\infty}\times(-\infty,0]\to\mathbb{R}^{d+1}$ where $X_{j}\to X_{\infty}$ smoothly on compact subsets of $M^{d}_{\infty}\times(-\infty,0]$. Also the ’flatness at low curvature region’ condition implies that
@@
This means $-\epsilon_{0}=\frac{\lambda_{1}+..+\lambda_{k}}{H}(p_{\infty},0)$ is a strictly negative spacetime interior minimum of $\frac{\lambda_{1}+..+\lambda_{k}}{H}$. In the case when $k=1$, this contradicts [5], proposition A.1, and when $k=d-1$, this contradicts lemma 3.2. Thus in any case, $-\epsilon_{0}\geq 0$, thus we have weak $k$ convexity when $k=1$ or $k=d-1$, thus proving theorem 1.4.
Remark 3.1
The extra ’flatness’ condition is only used to prevent the limit flow from being a static hyperplane. Also, note that the scaling $\phi(t_{j})$ is chosen to be able to extract a smooth limit. Therefore, as long as we can find some other conditions which ensures smooth convergence to a nonflat limit by possibly taking different scales, the argument holds without extra modification.

### v2

@@ 1 Introduction
Let $M^{d}$ be a smooth $d$-dimensional manifold. A smooth one-parameter family of hypersurfaces $\{M_{t}^{d}\}_{t\in I}$ is a smooth solution to mean curvature flow if there exists a smooth one parameter family of immersions $X_{t}:M^{d}\to\mathbb{R}^{d+1}$ which satisfies the equation
@@
A standard method of analyzing singularities is to use blowup arguments. As the solution approaches singular time, we ‘magnify’ at larger and larger scale near the point where singularities form, and take ‘the limit’ of the rescaled solutions. In the end, one is left with a solution to the mean curvature flow which is defined for all negative time. Such solutions are called ancient solutions. Therefore, the study of ancient solutions is important since they model how singularities form.
@@
Corollary 1.1 (Recovery of classical Harnack inequality).
@@
In the second half of the paper, we consider ancient solutions to mean curvature flow. In [6], Huisken and Sinestrari provided several characterizations of shrinking spheres for closed, convex ancient solution to mean curvature flow.
Theorem 1.2 (cf. [6], theorem 1.1).
@@
(2) $\liminf_{t\to-\infty}\min_{M^{d}\times\{t\}}\frac{\lambda_{1}}{H}>0$ i.e $\{M_{t}^{d}\}_{t\in(-\infty,0)}$ is uniformly strictly convex.
(3) $\limsup_{t\to-\infty}\frac{\textit{{diam}}(M_{t}^{d})}{\sqrt{-t}}<\infty$.
@@
i.e $\{M_{t}^{d}\}_{t\in(-\infty,0)}$ is uniformly strictly convex.
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}H\textit{{diam}}(M^{d}_{t})<\infty$
where $\textit{{diam}}(M^{d}_{t})$ denote the diameter of $M_{t}^{d}\subset\mathbb{R}^{d+1}$.
@@
for either $k=1$ or $k=d-1$. We also impose a ‘flatness at low curvature region’ condition, that is for every $\epsilon>0$, there exists $\delta>0$ so that whenever
@@
Acknowledgment
The author would like to thank his advisor Prof. Natasa Sesum for her various suggestions and comments to improve the results of this paper.
@@ 2 Generalized Harnack inequality
Proof of theorem 1.1.
For simplicity, we define
@@
$Z_{\epsilon,\hat{T}}(V)$ $=(A+\epsilon Hg)(V,V)+\frac{2}{1+\epsilon d}\nabla H\cdot V+\frac{1}{(1+\epsilon d)^{2}}\nabla_{t}H+\frac{H}{2t}+\frac{3\epsilon}{4(1+\epsilon d)}\phi(\hat{T})^{2}H+\frac{\epsilon}{(1+\epsilon d)^{2}}H^{3}$
@@
for all $V\in TM^{d}$, $0<t\leq\hat{T}<T$. Then theorem 1.1 follows by first letting $\epsilon\to\epsilon_{0}$, then $t=\hat{T}\in(0,T)$. Since $\{M^{d}_{t}\}_{t\in[0,T)}$ is a mean convex solution to mean curvature flow, by the strong maximum principle either $H>0$ for all $t\in(0,T)$ or $H=0$. The theorem trivially holds when $H=0$, thus WLOG we assume $H>0$ for $t>0$. Then for any $\epsilon>\epsilon_{0}$,
@@
For each $(p_{0},t_{0})\in M^{d}\times(0,\hat{T}]$, we let $(U,\{x_{a}\})$ to be a geodesic normal coordinates w.r.t $g(t_{0})$ centered at $p_{0}$ so that the coordinate vectors at $p_{0}$ are the eigenvectors of the second fundamental form i.e at $(p_{0},t_{0})$,
@@
$+(-4\nabla_{c}A_{ab}V_{b}-4\epsilon\nabla_{c}Hg_{ab}V_{b}-4\delta\nabla_{t}A_{ac}+4\delta HA_{ab}A_{bc})\nabla_{c}V_{a}+4\delta H\nabla_{a}A_{bc}A_{bc}V_{a}$
$+2\delta A_{ab}A_{bc}\nabla_{c}HV_{a}+4\delta^{2}HA_{ab}\nabla_{t}A_{ab}+2\delta^{2}A_{ab}\nabla_{a}H\nabla_{b}H-2\delta^{2}H^{2}A_{ab}A_{bc}A_{ca}$
$+2\epsilon\delta^{2}|A|^{2}H^{3}+g^{\prime}(t)H-6\epsilon\delta^{2}H|\nabla H|^{2}$
We compute each terms separately according to our choice of local extension.
@@
Combining all the terms, we obtain
@@
$(\nabla_{t}-\Delta)Z_{\epsilon,\hat{T}}(V)=$ $|A|^{2}Z_{\epsilon,\hat{T}}(V)-\frac{4f}{\delta}[(A_{ab}+\epsilon Hg_{ab})V_{a}V_{b}+2\delta\nabla_{a}HV_{a}+\delta^{2}\nabla_{t}H]$
$+(6\epsilon^{2}\delta H^{2}-4\epsilon f)\nabla_{a}HV_{a}+g^{\prime}(t)H-2f^{2}H-4f\epsilon\delta H^{3}-2\epsilon df^{2}H$
$=$ $|A|^{2}Z_{\epsilon,\hat{T}}(V)-\frac{4f}{\delta}[Z_{\epsilon,\hat{T}}(V)-g(t)H-\epsilon\delta^{2}H^{3}]$
$+(6\epsilon^{2}\delta H^{2}-4\epsilon f)\nabla_{a}HV_{a}+g^{\prime}(t)H-2f^{2}H-4f\epsilon\delta H^{3}-2\epsilon df^{2}H$
@@
and $(p_{0},t_{0})\in M^{d}\times(0,\hat{T}]$ and $\phi(\hat{T})=\sup_{M^{d}\times(0,\hat{T}]}H$. One can directly check that above two inequalities indeed hold, thus proving the desired differential inequality.
@@
for $V\in TM^{d}$, $t\in(0,\hat{T}]$, hence prove theorem 1.1. ∎
@@
Corollary 2.1 (Harnack inequality for ancient solutions).
@@
Proof.
For each fixed $V\in TM^{d}$, $t\in(-\infty,0)$, let $-\alpha<t<0$ and consider $\{M_{s}^{d}\}_{s\in[-\alpha,0)}$. Since $\phi(t)\geq\sup_{M^{d}\times(-\alpha,t]}H$, theorem 1.1 implies
@@
for all $V\in TM^{d}$, $t\in(-\alpha,0)$. Thus letting $\alpha\to\infty$, we obtain corollary 2.1. ∎
@@
Proof.
Let $\gamma:[t_{1},t_{2}]\to M^{d}$ be any smooth curve connecting $p_{1}$ and $p_{2}$. Dividing the Harnack inequality by $H$ and setting $V=\frac{1}{2(1+\epsilon_{0}d)}\gamma^{\prime}(t)$, we have
@@
thus proving corollary 2.2. ∎
@@ 3 Ancient solutions
$\limsup_{t\to-\infty}\phi(t)\textit{{diam}}(M_{t}^{d})<\infty\iff\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}H\textit{{diam}}(M_{t}^{d})<\infty$
Proof.
By definition of $\phi$,
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}H\textit{{diam}}(M_{t}^{d})\leq\limsup_{t\to-\infty}\phi(t)\textit{{diam}}(M_{t}^{d})$
@@
To prove the reverse direction, note that $t\to\textit{{diam}}(M_{t}^{d})=\max_{x,y\in M^{d}}|X_{t}(x)-X_{t}(y)|$ is nonincreasing in $t$. We give a brief sketch of why this is true. Since we are working with smooth solutions to the mean curvature flow, for each $t\in(-\infty,0)$
$\textit{{diam}}(M_{t}^{d})=\max_{x,y\in M^{d}}|X_{t}(x)-X_{t}(y)|>0$
@@
$\frac{d}{dt}\textit{{diam}}(M_{t}^{d})=\frac{\partial}{\partial t}|X_{t}(p)-X_{t}(q)|=\langle\frac{X_{t}(p)-X_{t}(q)}{\textit{{diam}}(M_{t}^{d})},\vec{H}(p,t)-\vec{H}(q,t)\rangle$
@@
$0\geq\langle\frac{X_{t}(p)-X_{t}(q)}{\textit{{diam}}(M_{t}^{d})},\vec{H}(p,t)-\vec{H}(q,t)\rangle+\frac{2d}{\textit{{diam}}(M_{t}^{d})}=\frac{d}{dt}\textit{{diam}}(M_{t}^{d})+\frac{2d}{\textit{{diam}}(M_{t}^{d})}$
which proves that $\textit{{diam}}(M_{t}^{d})$ is nonincreasing in $t$.
@@
$\limsup_{t\to-\infty}\max_{M^{d}\times\{t\}}H\textit{{diam}}(M_{t}^{d})\leq C<\infty$
@@
Since $\textit{{diam}}(M_{t}^{d})\leq\textit{{diam}}(M_{s}^{d})$, we have
$(\phi(t)-\epsilon)\textit{{diam}}(M_{t}^{d})\leq\max_{M^{d}\times\{s\}}H\textit{{diam}}(M_{s}^{d})\leq C+1$
@@
$\limsup_{t\to-\infty}\phi(t)\textit{{diam}}(M^{d}_{t})\leq C+1<\infty$
thus proving the lemma. ∎
@@
Proof.
Define a tensor
@@
$\lambda_{i}\lambda_{d}=\textit{{sec}}(e_{i},e_{d})=0$
hence $\lambda_{i}=0$. Here, $\textit{{sec}}(e_{i},e_{d})$ is the sectional curvature in directions $e_{i}$ and $e_{d}$. This together with the fact that $E(t)$ has rank $l>0$ implies that at each $(p,t)\in M^{d}\times(-\delta,0)$, the eigenvalues of $A_{(p,t)}$ are either 00 or $\lambda_{d}$, with the corresponding eigenspaces having dimension $d-l$ and $l$ respectively. Thus for any possible $l>0$, $A\geq 0$ in $M^{d}\times(-\delta,0)$. Then by continuity, $A\geq 0$ up to time $t=0$ as well, thus proving the first assertion.
@@
for any $q\in M^{d}$. ∎
We now prove theorem 1.3 and theorem 1.4.
Proof of theorem 1.3.
Trivially $(1)$ implies $(2)-(5)$. By [6], theorem 1.1, $(2)$ implies $(1)$. We claim that either $(4)$ or $(5)$ implies $(3)$. By the weak maximum principle, it is easy to see that for all sufficiently small $t$
@@
$\textit{{diam}}(M_{t}^{d})\leq 2C_{2}\sqrt{-t}+\textit{{diam}}(M_{T_{0}}^{d})\leq C_{3}\sqrt{-t}\leq\frac{C_{4}}{\phi(t)}$
@@
$\limsup_{t\to-\infty}\phi(t)\textit{{diam}}(M_{t}^{d})<\infty$
@@
is the smallest eigenvalue of $Hg-A$, we conclude that
@@
Then lemma 3.2 implies that $M_{\infty}^{d}$ is closed and $X_{\infty}(\cdot,0)$ is an isometric embedding into $\mathbb{R}^{d+1}$ with its image being a round sphere. In particular, the local smooth convergence at $t=0$ is actually global smooth convergence, hence the rescaled second fundamental form $\phi(t_{j})^{-1}A(\cdot,t_{j})$ converge uniformly smoothly to the second fundamental form of a round sphere of fixed radius as $j\to\infty$. Since a sphere is uniformly strictly convex, there exists $j_{0}$ so that for all $j\geq j_{0}$
@@
$\limsup_{t\to-\infty}\frac{\textit{{diam}}(M_{t}^{d})}{\sqrt{-t}}\leq C\limsup_{t\to-\infty}\phi(t)\textit{{diam}}(M_{t}^{d})<\infty$
thus by [6], theorem 1.1, $\{M_{t}^{d}\}_{t\in(-\infty,0)}$ is a family of shrinking round spheres. Therefore either (3), (4) or (5) implies (1), thus completing the proof of theorem 1.3. ∎
Proof of theorem 1.4.
Assume otherwise. Then
@@
defined for $t\in(-\infty,0]$. Then as in the proof of theorem 1.3, we can extract a limit flow $X_{\infty}:M^{d}_{\infty}\times(-\infty,0]\to\mathbb{R}^{d+1}$ where $X_{j}\to X_{\infty}$ smoothly on compact subsets of $M^{d}_{\infty}\times(-\infty,0]$. Also the ‘flatness at low curvature region’ condition implies that
@@
This means $-\epsilon_{0}=\frac{\lambda_{1}+..+\lambda_{k}}{H}(p_{\infty},0)$ is a strictly negative spacetime interior minimum of $\frac{\lambda_{1}+..+\lambda_{k}}{H}$. In the case when $k=1$, this contradicts [5], proposition A.1, and when $k=d-1$, this contradicts lemma 3.2. Thus in any case, $-\epsilon_{0}\geq 0$, thus we have weak $k$ convexity when $k=1$ or $k=d-1$, thus proving theorem 1.4. ∎
Remark 3.1.
The extra ‘flatness’ condition is only used to prevent the limit flow from being a static hyperplane. Also, note that the scaling factor $\phi(t_{j})$ is chosen to be able to extract a smooth limit. Therefore, as long as we can find some other conditions which ensures smooth convergence to a nonflat limit by possibly taking different scales, the argument holds without extra modification.